Can someone please help me with this problem. Thanks
A rectangular piece of tin measures 36 inches by 25 inches. A square measuring 4 inches is cut out of each corner and the sides are folded to form a box. What is the volume of the box?

Answer by rapaljer(4671)

If you cut 4 inches from each corner, you will reduce the dimensions of the base of the rectangular base by 8 inches, which would be 36-8 = 28 inches by 25-8 = 17 inches. The height of the box is of course 4 inches.

Volume = L*W*H
V= 28*17*4 = 1904 cubic inches.

R^2 at SCC